A gas sample at 0C is heated until the average velocity of its molecules doubles. What is its new temperature?
1
Expert's answer
2019-12-16T11:11:27-0500

Average velocity of gas molecules = (8RT/"\pi"M)1/2

where T being the temperature in kelvin , M is the molecular mass

we can say that velocity is directly proportional to the square root of T

then, V1/V2 =(T1/T2)1/2

as velocity gets doubled so V1/V2 = 1/2

so, 0.5= (273/T2)1/2

T2 = 273/0.25

=1092 K

In 0C =1092-273 = 819 0C